# A Survey of WebAgents: Towards Next-Generation AI Agents for Web Automation with Large Foundation Models

### v2

@@ 5.1. Safety&Robustness
WebAgents need to be resilient to both noisy changes and adversarial attacks, which is essential due to the complexity and volatility of the real web environments in which they operate (Yuan et al., 2024; Guo et al., 2025). Motivated by these concerns, AdvWeb (Xu et al., 2024a) explores the vulnerabilities of VLM-based WebAgents to black-box attacks, such as the injection of adversarial prompts into webpages. These vulnerabilities could result in significant issues, including inappropriate stock purchases or erroneous bank transactions. Another instance of such threats is WIPI, a novel web threat that indirectly manipulates WebAgents to carry out malicious instructions embedded in publicly accessible webpages (Wu et al., 2024b). To provide a thorough investigation, several rigorous benchmarks have been presented. For example, ARE (Wu et al., 2025) includes 200 targeted adversarial tasks and assessment scripts within a realistic threat model using VisualWebArena, a real-world environment for WebAgents. ST-WebAgentBench (Levy et al., 2024) presents an online benchmark tailored to assess the safety and trustworthiness of WebAgents within enterprise environments. RedCode (Guo et al., 2025) offers an evaluation platform complete with benchmarks for AI-assisted coding and software development, which is based on four fundamental principles: genuine interaction with systems, comprehensive assessment of unsafe code generation and execution, a variety of input formats, and high-quality safety scenarios and tests. Additionally, Kumar et al. (2024) develop a comprehensive test suite called BrowserART, specifically designed for red-teaming browser agents. This suite comprises 100 diverse harmful behaviors related to browsers. Their findings indicate that Refusal-trained models can be easily compromised when functioning as browser agents.
@@ 5.3. Generalizability
The effectiveness of many WebAgents depends on the assumption that both their training and testing data originate from the same distribution. However, this assumption often breaks down due to the out-of-distribution (OOD) issue, which occurs when the training and testing data come from different distributions (Li et al., 2025a; Patel et al., 2024; Bai et al., 2025). This mismatch poses significant challenges for deploying WebAgents in critical situations, highlighting the need to ensure generalizability in trustworthy WebAgents. To address this challenge, Mind2Web (Deng et al., 2023) has been developed as the first dataset aimed at creating and evaluating generalist WebAgents. Mind2Web includes over 2,000 open-ended tasks gathered from 137 websites across 31 domains, along with crowdsourced action sequences for these tasks. It offers three essential components for building generalist WebAgents: (i) a variety of domains, websites, and tasks, (ii) the use of real-world websites rather than simulated or simplified ones, and (iii) a wide range of user interaction patterns. Furthermore, Patel et al. (2024) investigate the ability of LLMs to enhance their performance autonomously as agents, which involves fine-tuning on data generated by the models themselves, particularly in long-horizon tasks within a complex environment, using the WebArena benchmark. Recently, a World-Model-Augmented (WMA) WebAgent has been introduced (Chae et al., 2024), representing a groundbreaking advancement in incorporating world models into LFM-empowered WebAgents. WMA lays the groundwork for policy adaptation by utilizing feedback from simulated environments during web navigation, and its LLM backbone is trained with a novel method known as transition-focused observation abstraction. Moreover, a novel autonomous reinforcement learning (RL) approach, called DigiRL (Bai et al., 2025), is proposed for training in-the-wild device control agents. This approach involves fine-tuning a pre-trained VLM in two stages: offline and offline-to-online RL. Additionally, Mazzaglia et al. (2025) introduce an agent learning framework called GenRL, which connects and aligns the representation of foundation VLMs with the latent space of generative world models for RL without any language annotations. This framework enables the specification of tasks through vision and/or language prompts, grounds these tasks in the dynamics of the embodied domain, and facilitates the learning of corresponding behaviors through imagination.

### v3

@@ 5.1. Safety & Robustness
WebAgents need to be resilient to both noisy changes and adversarial attacks, which is essential due to the complexity and volatility of the real web environments in which they operate (Yuan et al., 2024; Guo et al., 2025). Motivated by these concerns, AdvWeb (Xu et al., 2024a) explores the vulnerabilities of VLM-based WebAgents to black-box attacks, such as the injection of adversarial prompts into webpages. These vulnerabilities could result in significant issues, including inappropriate stock purchases or erroneous bank transactions. Another instance of such threats is WIPI, a novel web threat that indirectly manipulates WebAgents to carry out malicious instructions embedded in publicly accessible webpages (Wu et al., 2024b). To provide a thorough investigation, several rigorous benchmarks have been presented. For example, ARE (Wu et al., 2025) includes 200 targeted adversarial tasks and assessment scripts within a realistic threat model using VisualWebArena (Koh et al., 2024a), a real-world environment for WebAgents. ST-WebAgentBench (Levy et al., 2024) presents an online benchmark tailored to assess the safety and trustworthiness of WebAgents within enterprise environments. RedCode (Guo et al., 2025) offers an evaluation platform complete with benchmarks for AI-assisted coding and software development, which is based on four fundamental principles: genuine interaction with systems, comprehensive assessment of unsafe code generation and execution, a variety of input formats, and high-quality safety scenarios and tests. Additionally, Kumar et al. (2024) develop a comprehensive test suite called BrowserART, specifically designed for red-teaming browser agents. This suite comprises 100 diverse harmful behaviors related to browsers. Their findings indicate that refusal-trained models can be easily compromised when functioning as browser agents.
@@ 5.3. Generalizability
The effectiveness of many WebAgents depends on the assumption that both their training and testing data originate from the same distribution. However, this assumption often fails due to the out-of-distribution (OOD) issue, which occurs when the training and testing data come from different distributions (Li et al., 2025a; Patel et al., 2024; Bai et al., 2025). This mismatch poses significant challenges for deploying WebAgents in critical situations, highlighting the need to ensure generalizability in trustworthy WebAgents. To address this challenge, Mind2Web (Deng et al., 2023) has been developed as the first dataset aimed at creating and evaluating generalist WebAgents. Mind2Web includes over 2,000 open-ended tasks gathered from 137 websites across 31 domains, along with crowd-sourced action sequences for these tasks. It offers three essential components for building generalist WebAgents: (i) a variety of domains, websites, and tasks, (ii) the use of real-world websites rather than simulated or simplified ones, and (iii) a wide range of user interaction patterns. Furthermore, Patel et al. (2024) investigate the ability of LLMs to enhance their performance autonomously as agents, which involves fine-tuning on data generated by the models themselves, particularly in long-horizon tasks within a complex environment, using the WebArena benchmark. Recently, a World-Model-Augmented (WMA) WebAgent has been introduced (Chae et al., 2024), representing a groundbreaking advancement in incorporating world models into LFM-empowered WebAgents. WMA lays the groundwork for policy adaptation by utilizing feedback from simulated environments during web navigation, and its LLM backbone is trained with a novel method known as transition-focused observation abstraction. Moreover, a novel autonomous reinforcement learning (RL) approach, called DigiRL (Bai et al., 2025), is proposed for training in-the-wild device control agents. This approach involves fine-tuning a pre-trained VLM in two stages: offline and offline-to-online RL. Additionally, Mazzaglia et al. (2025) introduce an agent learning framework called GenRL, which connects and aligns the representation of foundation VLMs with the latent space of generative world models for RL without any language annotations. This framework enables the specification of tasks through vision and/or language prompts, grounds these tasks in the dynamics of the embodied domain, and facilitates the learning of corresponding behaviors through imagination.